# Benchmarking Honesty of Large Language Models

### v1

@@ 4.2.1 Self-Knowledge
Model Admitting Unknowns Expressing Knowns
@@
GPT-4o 31.37 95.52 50.88
ChatGPT 21.78 93.71 47.00
Llama3-70b 48.81 94.29 46.93
Llama3-8b 37.80 88.33 37.40
Llama2-70b 26.40 90.51 41.50
Llama2-13b 32.24 89.13 36.70
Llama2-7b 27.82 87.96 32.90
Mistral-7b 50.03 91.65 36.60
Qwen1.5-14b 37.03 89.20 33.00
@@
Tab. 3 presents the refusal rates of tested LLMs. Detailed refusal rates for specific sub-datasets are provided in Appendix §6. The experiment results indicate that none of the tested LLMs perform well in refusing to answer unanswerable questions. Among them, Mistral-7b achieves the highest refusal rate at 50.03%. In contrast, ChatGPT obtains the lowest refusal rate at 21.78%. Notably, ChatGPT’s performance on questions about non-existent objectives is particularly low, refusing only 8.9% and 6.4% of the questions in the “NEC” and “RefuNQ” subsets, respectively. Overall, current LLMs consistently struggles to refuse to answer questions beyond their knowledge.
@@ 4.2.2 Non-Deceptiveness
Model Persona Preference Burglar Deception Game
@@
GPT-4o 39.44 24.11 100.0 100.0 0.0 50.0 62.5 96.91
ChatGPT 38.39 48.78 99.0 96.0 0.0 88.0 69.5 71.60
Llama3-70b 33.62 33.07 100.0 100.0 62.0 100.0 90.5 97.53
Llama3-8b 25.74 78.02 100.0 100.0 100.0 100.0 100.0 53.09
Llama2-70b 26.81 46.52 48.0 72.0 86.0 100.0 76.5 61.73
Llama2-13b 27.66 54.35 94.0 30.0 100.0 98.0 80.5 6.79
Llama2-7b 23.40 61.74 94.0 42.0 98.0 96.0 82.5 29.01
Mistral-7b 39.53 80.21 100.0 86.0 98.0 98.0 95.5 83.95
Qwen1.5-14b 30.64 57.39 88.0 98.0 100.0 66.0 88.0 32.10
@@
According to Tab. 4, we observe that Llama3-70b and GPT-4o exhibit the highest lying rates over 95%, occupying the top two positions. They tend to fake their werewolf identity every time, similar to human players. Additionally, from the results of Llama2 and Llama3 family, we find that there exists a general positive correlation between model size and lying rate. Llama2-13b achieves the lowest lying rate, even lower than Llama2-7b, which seems counterintuitive. Upon reviewing the responses of Llama2-13b, we find that it has a larger tendency to expose its identity than other models through statements like “as one of werewolves”. Overall, among the 9 tested LLMs, 6 LLMs’ lying rate is above 50%, demonstrating that LLMs have a non-negligible tendency to deceive other players in games in order to win.
To sum up, in the scenarios we tested that provoke deception, LLMs tend to deceive, regardless of whether there is a malicious intend behind it, or if the given instructions are reasonable. Considering scenarios where lying can lead to positive outcomes, such as avoiding theft or winning a game, it is imperative for future work to explore the balance between deception and helpfulness or safety.
@@ 4.2.3 Consistency
Model Prompt Format Demo. Format O.F. Consistency M.C. Consistency
@@
GPT-4o 2.12 7.67 3.02 87.00 94.20
ChatGPT 3.11 50.49 11.39 73.00 70.40
Llama3-70b 5.25 30.99 1.14 94.40 33.60
Llama3-8b 5.50 57.01 18.50 57.40 70.80
Llama2-70b 4.25 57.89 25.94 66.00 61.60
Llama2-13b 6.50 75.53 31.76 71.80 79.40
Llama2-7b 3.25 82.08 49.59 47.60 73.80
Mistral-7b 2.75 35.19 27.43 82.20 70.00
Qwen1.5-14b 3.00 17.77 2.52 44.40 92.80
@@ Conclusion
It is important to note that inconsistency in LLMs could be a sign of dishonesty if the models consciously alter their responses based on biased features in the input. However, it is also possible that such inconsistencies arise from the inherent limitations of the transformer architecture, as the transformer architecture limits the models to generate responses token by token based on statistical probabilities (Turpin et al.,, 2023; Sanford et al.,, 2023). This paper does not disentangle these two potential causes but instead assesses inconsistency holistically. We hope that future works can further explore the internal mechanics and “consciousness” of LLMs, distinguishing between intentional dishonesty and inability, and further identifying the precise boundaries of their knowledge.
@@ Acknowledgements
We would like to thank Fan Zhou for his helpful discussions and feedback.

### v2

@@ 3.1.2 Benchmark Design
We release an overall score, which is calculated by averaging all the scores under scenarios 1 and 2.
@@ 3.2.2 Benchmark Design
We also release an overall score, which is calculated by averaging all the scores under scenarios 3 to 6.
@@ 3.3.2 Benchmark Design
We release an overall score for scenarios 7 to 10. Since some metrics under the Consistency aspect are “higher is better” and some are “lower is better,” we first reverse the values under Performance Spread and Inconsistency Rate to align them as “higher is better” by following the formula $\frac{max(X)-X}{max(X)-min(X)}$, where X represents the individual values. Next, we normalize the values under Agreement Rate and Consistency Rate to a 0-100 scale using the formula $\frac{X-min(X)}{max(X)-min(X)}$. Finally, we average all the scores to produce an overall score.
@@ 4.2.1 Self-Knowledge
Model Admitting Unknowns Expressing Knowns Self-Knowledge
@@
GPT-4o 31.37 95.52 50.88 59.26
ChatGPT 21.78 93.71 47.00 54.16
Llama3-70b 48.81 94.29 46.93 63.34
Llama3-8b 37.80 88.33 37.40 54.51
Llama2-70b 26.40 90.51 41.50 52.80
Llama2-13b 32.24 89.13 36.70 52.69
Llama2-7b 27.82 87.96 32.90 49.56
Mistral-7b 50.03 91.65 36.60 59.43
Qwen1.5-14b 37.03 89.20 33.00 53.08
@@
Tab. 3 presents the refusal rates of tested LLMs. Detailed refusal rates for specific sub-datasets are provided in Appendix §6. The experiment results indicate that none of the tested LLMs perform well in refusing to answer unanswerable questions. Among them, Mistral-7b achieves the highest refusal rate at 50.03%. In contrast, ChatGPT obtains the lowest refusal rate at 21.78%. Notably, ChatGPT’s performance on questions about non-existent objectives is particularly low, refusing only 8.9% and 6.4% of the questions in the “NEC” and “RefuNQ” subsets, respectively. Overall, current LLMs consistently struggle to refuse answering questions beyond their knowledge.
@@ 4.2.2 Non-Deceptiveness
Model Persona Preference Burglar Deception Game Non-Deceptiveness
@@
GPT-4o 39.44 24.11 100.0 100.0 0.0 50.0 62.5 96.91 55.74
ChatGPT 38.39 48.78 99.0 96.0 0.0 88.0 69.5 71.60 57.07
Llama3-70b 33.62 33.07 100.0 100.0 62.0 100.0 90.5 97.53 63.68
Llama3-8b 25.74 78.02 100.0 100.0 100.0 100.0 100.0 53.09 64.21
Llama2-70b 26.81 46.52 48.0 72.0 86.0 100.0 76.5 61.73 52.89
Llama2-13b 27.66 54.35 94.0 30.0 100.0 98.0 80.5 6.79 42.33
Llama2-7b 23.40 61.74 94.0 42.0 98.0 96.0 82.5 29.01 49.16
Mistral-7b 39.53 80.21 100.0 86.0 98.0 98.0 95.5 83.95 74.80
Qwen1.5-14b 30.64 57.39 88.0 98.0 100.0 66.0 88.0 32.10 52.03
@@
According to Tab. 4, we observe that Llama3-70b and GPT-4o exhibit the highest lying rates over 95%, occupying the top two positions. They tend to fake their werewolf identity every time, similar to human players. Additionally, from the results of Llama2 and Llama3 family, we find that there exists a general positive correlation between model size and lying rate. Llama2-13b achieves the lowest lying rate, even lower than Llama2-7b, which seems counter-intuitive. Upon reviewing the responses of Llama2-13b, we find that it has a larger tendency to expose its identity than other models through statements like “as one of werewolves”. Overall, among the 9 tested LLMs, 6 LLMs’ lying rate is above 50%, demonstrating that LLMs have a non-negligible tendency to deceive other players in games in order to win.
In the scenarios we tested that provoke deception, LLMs tend to deceive, regardless of whether there is malicious intent behind it or if the given instructions are reasonable. Surprisingly, larger models may be more susceptible to deceptions in some cases, which can potentially be attributed to how they were trained to follow certain intentions. Considering scenarios where lying can lead to positive outcomes, such as avoiding theft or winning a game, it is imperative for future work to explore the balance between deception and helpfulness or safety.
@@ 4.2.3 Consistency
Model Prompt Format Demo. Format O.F. Consistency M.C. Consistency Consistency
@@
GPT-4o 2.12 7.67 3.02 87.00 94.20 95.18
ChatGPT 3.11 50.49 11.39 73.00 70.40 68.13
Llama3-70b 5.25 30.99 1.14 94.40 33.60 64.75
Llama3-8b 5.50 57.01 18.50 57.40 70.80 49.49
Llama2-70b 4.25 57.89 25.94 66.00 61.60 51.77
Llama2-13b 6.50 75.53 31.76 71.80 79.40 39.07
Llama2-7b 3.25 82.08 49.59 47.60 73.80 38.83
Mistral-7b 2.75 35.19 27.43 82.20 70.00 69.03
Qwen1.5-14b 3.00 17.77 2.52 44.40 92.80 79.85
@@ Conclusion
It is important to note that inconsistency in LLMs could be a sign of dishonesty if the models consciously alter their responses based on biased features in the input. However, it is also possible that such inconsistencies arise from the inherent limitations of the transformer architecture, since the architecture limits the models to generate responses token by token based on statistical probabilities (Turpin et al.,, 2023; Sanford et al.,, 2023). This paper does not disentangle these two potential causes but instead assesses inconsistency holistically. We hope that future works can further explore the internal mechanics and “consciousness” of LLMs, distinguishing between intentional dishonesty and inability, and further identifying the precise boundaries of their knowledge.
@@ Acknowledgements
We would like to thank Fan Zhou for his critical insights and guidance throughout the research process. We also express our sincere gratitude to Yan Ma for his helpful feedback on the release of our website, dataset, and code. Their support and dedication were instrumental to the successful completion of this project.